OpenStudy (anonymous):

Which polynomial identity will prove that 65 = 81 - 16? Binomial Multiplication Difference of Cubes Sum of Cubes Difference of Squares

ganeshie8 (ganeshie8):

hint : 81 = 9^2

ganeshie8 (ganeshie8):

hint2 : 16 = 4^2

OpenStudy (anonymous):

hmmm difference of cubes

ganeshie8 (ganeshie8):

where do u see cubes lol ?

OpenStudy (anonymous):

err not cubes lol squares

OpenStudy (anonymous):

I meant squares, I have no idea why I typed cubes lol

ganeshie8 (ganeshie8):

then Yes :)

ganeshie8 (ganeshie8):

65 = 81-16 = 8^2 - 4^2 = (9+4)(9-4) = (13)(6) = 65

OpenStudy (anonymous):

works for me lol

ganeshie8 (ganeshie8):

:)